OpenStudy (anonymous):

Venus is about 10^8 kilometers from the Sun. Saturn is about 10^9 kilometers from the Sun. About how many times farther from the Sun is Saturn than Venus?

OpenStudy (anonymous):

Hello, let me know if you have question to better explain myself. What is 10^n all about, Explanation: 10^9, for example, means to multiply 10 nine times. Looks like: 10*10*10*10*10*10*10*10*10=1,000,000,000 How?: 10*10=100 now times 100 with another 10=1,000 and so on for 7 times more for 10^9. You would multiply 10 eight times for 10^8. It all depends on the exponent number ( ^n ). Saturn is 10^9 km from the Sun. Saturn is farther from the sun compared to Venus that is 10^8 km from the sun. To know how farther is Saturn than Venus from the sun, subtract both distances. Saturn 10^9 - Venus 10^8 = 10^1 To answer the question, how many times more is Saturn farther than Venus: Saturn is 1 time more farther from the sun than Venus.